OpenStudy (anonymous):

What would the fraction 28/9 be if you turned it into a mixed number?

OpenStudy (addicted2anime):

3 2/9

OpenStudy (mathstudent55):

9 * 3 = 27, so 28/9 = 3 with remainder 1 \(\dfrac{28}{9} = 3 \dfrac{1}{9} \)

OpenStudy (damonte212):

3 2/9